(The following information applies to the questions displayed below.] The following events apply to Gulf Seafood for the Yeara. Record the above transactions in a horizontal statements model. (In the Statement of Cash Flows column, use the initials (b. What amount of depreciation expense would Gulf Seafood report on the Year 1 income statement? Depreciation expensec. What amount of accumulated depreciation would Gulf Seafood report on the December 31, Year 2, balance sheet? Accumulated dd. Would the cash flow from operating activities be affected by depreciation in Year 1? Yes No
Answer #1
GULF SEAFOOD
Horizontal Statements Model
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Stockholder's Equity Revenue - Expense = Net Income
Cash + Equipment (BV) = Common Stock + Retained Earnings
1 $18000 + $0 = $18000 + $0 $0 - $0 = $0 $18000
2 ($13600) + $13600 = $0 + $0 $0 - $0 = $0 ($13600)
3 $23000 + $0 = $0 + $23000 $23000 - $0 = $23000 $23000
4 ($11500) + 0 = $0 + ($11500) $0 - $11500 = ($11500) ($11500)
5 $0 + ($2900) = $0 + ($2900) $0 - $2900 = ($2900) $0
Balance $15900 + $10700 = $18000 + $8600 $23000 - $14400 = $8600 $15900

b) Depreciation for the first year

Cost. of asset =$13600

Salvage value =$2000

Useful life = 4 years

Depreciation = ( 13600-2000) ÷ 4

= $ 2900

c) Accumulated depreciation

Accumulated depreciation on 2 year = 2900 × 2

= $ 5800

d) No

Depreciation is a non cash expense . So it will not affect the cash flow.

Note : Accumulated depreciation for first year is shown as a negative figure in equipment column in the table as column for it is not provided in the table.

Please provide feedback if it was helpful. Feel free to comment if you have any doubts.